mu=40, sigma=10. If random samples of size n = 25 are selected from this population, what proportion of sample means will have values between 38 and 42?
Answer by stanbon(75887)

mu=40, sigma=10. If random samples of size n = 25 are selected from this population, what proportion of sample means will have values between 38 and 42?
-----------------
z(38) = (38-40)/10 = -2/10 = -0.2
---
z(42) = (42-40)/10 = +0.2
--------
P(38< x <42) = P(-0.2< z <0.2) = normalcdf(-0.2,0.2) = 0.1585
-------------------
cheers,
Stan H.
====================